31.5 g NH4Cl is dissolved in 500.ml H2O initially at 23c . Determine the final temperature of the resulting solution


1
Expert's answer
2021-10-13T04:47:02-0400

V2=0.5l*T2-23/10

V2=31.5/53.5*22.4=13.19l

13.2*10/0.5=T2-23

T2-23=264

T2=287